OpenStudy (anonymous):

Order of Operations-Evaluate this Expression-72-3Over(2+2)^2

OpenStudy (anonymous):

PEMDAS P= parentheses E= exponents M/D= Multiplication/ Divition A/S= Addition/ Subtraction do the problem in that order :)

OpenStudy (anonymous):

hope that helps!!

OpenStudy (anonymous):

Is it 69 over 16?

OpenStudy (anonymous):

show me your work

OpenStudy (anonymous):

72-3=69 (2+2)^2=4^2=16

OpenStudy (anonymous):

correct